OpenStudy (anonymous):

The ratio of the number of cookies to the number of cupcakes in a box is 5:4. Which table shows the possible amounts of cookies and cupcakes in a box, given the ratio? Cookies Cupcakes 5 4 10 8 20 16 40 32 Cookies Cupcakes 5 4 10 9 20 19 40 39 Cookies Cupcakes 5 4 6 5 7 6 8 7 Cookies Cupcakes 5 4 7 6 9 8 11 10

OpenStudy (anonymous):

@DIDITHEDAD

OpenStudy (lukeskylicker):

This is a very simple one. The ration means that for every 5 cookies in the box there will be 4 cupcakes. So simply look for the table where this rule applies.

OpenStudy (anonymous):

Okay, divide all the numbers on the left by 5, and divide all the numbers on the right by 4. If they ALL come out to a normal number, meaning no decimal, you have your answer. :) So for A... Cookies Cupcakes 5 4 10 8 20 16 40 32 5/5 = ? 10/5 = ? 20/5 = ? 40/5 = ? 4/4 = ? 8/4 = ? 16/4 = ? 32/4 = ?

OpenStudy (anonymous):

ummm'

OpenStudy (anonymous):

@MissSmartiez

OpenStudy (anonymous):

@MissSmartiez

OpenStudy (anonymous):

@pink33

OpenStudy (anonymous):

@AloneS

OpenStudy (anonymous):

@MissSmartiez

OpenStudy (anonymous):

help plz

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

@TheSmartOne